OpenStudy (anonymous):

what is a possible first step in solving for x in the equation -3x-5+2x=6? a. subtract 6 both sides b. divide both sides by 6 c. divide both sides by -5 d. combine like terms

OpenStudy (anonymous):

simplify....that is, -3x + 2x = -x then the eqn becomes -x-5=6 then ad 5 to both sides to get, -x = 11 Ans. x=-11

OpenStudy (radar):

Note that it asking for a "possible" first step. I would combine similar terms first, getting: -x-5=6 Then I would add 5 to both sides getting -x=11 then multiply by -1 to change signs x=-11